# Vortices and Factorization

### v1

@@ 13 Conclusions and Open Problems
In the present paper we reviewed developments in application of factorization methods to the theory of point vortex patterns that took place during several last decades. In the $\Lambda=1$ case, these development have connections with the theory of KdV hierarchy of integrable PDEs and the bi-spectral problem. Some connections with integrable hierarchies have been also found in the $\Lambda=2$ case [17]. However, this case seems not to have as many exceptional properties: Absence of trigonometric counterparts of the main $\Lambda=2$ sequences (24,25) makes us think that this should be so.
In more detail: We took the trigonometric analog of the bilinear equation (8)
$\tau_{n-1}^{\prime\prime}\tau_{n}-2\gamma_{n}\tau_{n-1}^{\prime}\tau_{n}^{\prime}+\gamma_{n}^{2}\tau_{n-1}\tau_{n}^{\prime\prime}+(d_{n-1}-\gamma_{n}d_{n})^{2}\tau_{n-1}\tau_{n}=0,$ (102)
where $d_{n}=\deg\tau_{n}$, $\gamma_{n-1}\gamma_{n}=1$, $\gamma_{n}\in\{\Lambda^{-1},\Lambda\}$. In the $\Lambda=1$ case, sequence of its solutions, which is the trigonometric analog of the main sequence (Adler-Moser polynomials), is given by (77,78) with $k_{n}=n$. Here, the highest symbol of the $n$th Adler-Moser polynomial and its trigonometric counterpart is of degree $n(n+1)/2$, i.e. in the trigonometric case $\tau_{n}=\xi^{n(n+1)/2}+\alpha_{1}\xi^{n(n+1)/2-2}+\dots+\alpha_{n}\xi^{-n(n+1)/2}$, where $\xi=\exp({\rm i}z)$. For both types of polynomials a new free parameter appears at each step of the factorization chain.
However, when $\Lambda=2$, non-terminating solutions of the chain (102) with $\tau_{n}=\xi^{D_{n}}+\dots$, are (up to multiplication by constants and shift of $z$) only $\sin^{D_{n}}z$, where $D_{n}$ are degrees of the main sequences (24,25) [^18] We recall that, according to (63), for the main sequences (24,25) $D_{i+1}-2\gamma_{i}D_{i}+D_{i-1}=1$, $i\in{\mathbb{Z}}$ and $D_{0}=D_{1}=0$, $\gamma_{0}=2$.. We searched for other configurations imposing condition that $\tau_{i}$ do not have common or multiple roots, i.e. for configurations of two distinct species of vortices. By direct solution of (102), we found only short terminating sequences, for example
@@
where $\gamma_{0}=2$. The above example can be continued one more step if we impose condition $t_{1}=t_{2}^{2}/5$. Then we will get another terminating sequence with $\tau_{1}=\tau_{1}(z,t_{2})$, $\tau_{2}=\tau_{2}(z,t_{2})$ and $\tau_{3}=\tau_{3}(z,t_{2},t_{3})$ , where $\deg\tau_{3}=14$.
Without going into much detail we just note that in such type of sequences $\tau_{1}$ is a solution of an ODE that is reducible to the hypergeometric form. Then we can apply KWCC transformations (59) to $\psi_{i}=\tau_{i}^{\gamma_{i}}/\tau_{i-1}$, $i=1,2$, to continue the sequence. The transformation can be applied at most twice. Terminating configurations obtained in this way can depend on no more than two non-trivial free parameters (they depend on two parameters when $\deg\tau_{1}=6,8,10,\dots$). Here, a natural question arises: If the above procedure generates all possible static vortex streets configurations for two species of vortices with $Q_{i}\in\{-1,2\}$
Concerning the $\Lambda=2$ case, we also note that methods displayed in this review are inapplicable to both the translating streets and to the translating patterns composed of finite numbers of vortices when ${\cal Q}_{i}\in\{-1,2\}$.
Regarding the static patterns formed by finite number of vortices, we may ask the question: Whether the main sequences (23), (24), (25) for the $\Lambda=1$ and $\Lambda=2$ cases are the only possible non-trivial non-terminating sequences of the KWCC transforms. Are there non-trivial sequences (59) with $\gamma_{i}$ different from (60,61) that do not terminate?

### v2

@@ 13 Conclusions and Open Problems
In the present paper we reviewed developments in application of factorization methods to the theory of point vortex patterns that took place during several last decades. In the $\Lambda=1$ case, these have connections with the theory of KdV hierarchy of integrable PDEs and the bi-spectral problem. In the $\Lambda=2$ case, connections with integrable hierarchies have been found by Demina and Kudryashov [15] and further developed by us in [28]. In difference from the KdV hierarchy, the Lax operators of these hierarchies are differential operators of the third order  [^18] We recall that in $\Lambda=1$ case the Lax operator is of the second order: Lax operator of the KdV hierarchy is a Schrodinger operator.. Indeed, in the $\Lambda=2$ case the bilinear equation (8) for $q_{n}$ and $p_{n}$ can be rewritten in the form of the Schrodinger equation at zero energy level
$\left(-\partial_{z}^{2}+u\right)\phi=0,\quad u=-6(\log q_{n})^{\prime\prime},$ (103)
where $\phi=p_{n}/q_{n}^{2}$ or $\phi=p_{n-1}/q_{n}^{2}$. According to (24), we can replace $p_{n}$ with $q_{n+1}^{\prime}q_{n}-q_{n+1}q_{n}^{\prime}$ and get the following third-order differential equation for $\psi(z)=q_{n\pm 1}(z)/q_{n}(z)$
$L\psi=0,\quad L=\partial_{z}^{3}-u\partial_{z}.$
The third-order operator $L$ is a Lax operator of the Sawada-Kotera hierarchy and $q_{n}$ is its polynomial $\tau$-function. Such polynomials can be constructed with help of the Darboux transformations at “zero level” [28]
[…]
where primitives include arbitrary constants of integration. Corresponding third-order operators can be obtained by applications of sequence of the above transformations to $\partial_{z}^{3}$, while $q_{n}$ can be expressed in form of a Pfaffian [20].
Here we just note that, since $L$ is not a Schrodinger operator, the Baker-Akhiezer function $\psi(k,z)=P_{n}(k,z)/q_{n}(z)e^{kz}$ of the Sawada-Kotera hierarchy
$L\psi(k,z)=k^{3}\psi(k,z)$
does not correspond to stationary configurations when $k\not=0$. Also, soliton solutions of this hierarchy do not correspond to stationary configurations on the cylinder. Only when $k=0$, the above eigenvalue problem reduces to solution of the Schrodinger equation (103) with $\phi(z)=\lim_{k\to 0}k^{N_{n}}\frac{\partial\psi(k,z)}{\partial z}$ corresponding to static configurations. Here factor $k^{N_{n}}$, with suitably chosen $N_{n}$, insures finiteness of the limit.
In difference from the $\Lambda=1$ case, the $\Lambda=2$ case seems not to have as many exceptional properties: In particular, the fact that the soliton Baker-Akhiezer functions of the Sawada-Kotera hierarchy are not eigenfunctions of the Schrodinger operators suggests absence of non-trivial non-terminating sequences on the cylinder in the $\Lambda=2$ case.
We searched for sequences of configuration on the cylinder by direct solution of the trigonometric analog of the bilinear equation (8)
$\tau_{n-1}^{\prime\prime}\tau_{n}-2\gamma_{n}\tau_{n-1}^{\prime}\tau_{n}^{\prime}+\gamma_{n}^{2}\tau_{n-1}\tau_{n}^{\prime\prime}+(d_{n-1}-\gamma_{n}d_{n})^{2}\tau_{n-1}\tau_{n}=0,\quad\gamma_{n}\gamma_{n+1}=1,$ (104)
and in the $\Lambda=2$ case we found only short terminating sequences, for example
@@
where $\xi=e^{{\rm i}z}$ and $\gamma_{0}=2$. The above example can be continued one more step if we impose condition $t_{1}=t_{2}^{2}/5$. Then we will get another terminating sequence with $\tau_{1}=\tau_{1}(\xi,t_{2})$, $\tau_{2}=\tau_{2}(\xi,t_{2})$ and $\tau_{3}=\tau_{3}(\xi,t_{2},t_{3})$ , where $\deg\tau_{3}=14$.
Without going into much detail we just note that in such type of sequences $\tau_{1}$ is a solution of an ODE that is reducible to the hypergeometric form. Then we can apply KWCC transformations (60) to $\psi_{i}=\tau_{i}^{\gamma_{i}}/\tau_{i-1}$, $i=1,2$, to continue the sequence. The transformation can be applied at most twice. Terminating configurations obtained in this way can depend on no more than two non-trivial free parameters (they depend on two parameters when $\deg\tau_{1}=6,8,10,\dots$). Here, a natural question arises: If the above procedure generates all possible static vortex streets patterns of two distinct species of vortices with $Q_{i}\in\{-1,2\}$
Regarding the static patterns formed by finite number of vortices, we may ask the question: Whether the main sequences (23), (25), (26) for the $\Lambda=1$ and $\Lambda=2$ cases are the only possible non-trivial non-terminating sequences of the KWCC transforms. Are there non-trivial sequences (60) with $\gamma_{i}$ different from (61,62) that do not terminate?